an aluminum plate is made according to the following dimensions. The outer radius is 10 inches and the inner radius is 3 inches. Use p=3.14 as an approximate value. (a) how many square inches in area is the aluminum plate (the shaded region)? (b) If the cost to construct the plate is &7.50 per square inch, what is the construction cost?
Answer by checkley77(12844)

area=pir^2
area=3.14*5^2
area=3.14*25
area=78.5 in^2 for the total area of the 10 inch plate.
area=3.14*3^2
area=3.14*9
area=28.26 in^2 for the inner circle.
78.5-28.26=50.24 is the area of the 10 in. plate with a 3 in. hole.